OpenStudy (anonymous):

Multiply: 1/-7*5*-4/8

OpenStudy (ranga):

Is it \[\Large \frac{ 1 }{ -7 } \times 5 \times \frac{ -4 }{ 8 }\]?

OpenStudy (anonymous):

Yes

OpenStudy (ranga):

\[\Large \frac{ 1 }{ -7 } \times 5 \times \frac{ -4 }{ 8 } = \frac{ -20 }{ -56 } = \frac{ 5 }{ 14 }\]

OpenStudy (anonymous):

Thanks!

OpenStudy (ranga):

Welcome!